a single lane 12ft wide goes through a semicircular tunnel with radius 9ft. how high is the tunnel at the edge of each lane? (round off to 2 decimal places)
Answer by Boreal(15235)

The equation of the circle using the tunnel height as the y-axis and the width of the lane as the x-axis, and the origin in the middle is x^2+y^2=81
At the edge of the lane is at (6, 0) or (-6, 0)
so 36+y^2=81
y=sqrt(45) feet tall or 3 sqrt(5)=6.71 feet tall.